# Object-centric proto-symbolic behavioural reasoning from pixels

### v2

@@ Abstract
[inst2]organization=Cajal International Neuroscience Center, Spanish National Research Council, city=Madrid, country=Spain
@@ 1 Introduction
A one-year old infant, before language is expressed [1], learns the efferent-afferent patterns of sensory stimulation and motor commands to a neural representation of the environment. The pathway from the sensorium, to abstract thought, and back to the minutiae of the sensorimotor domain defines the feats of higher-order cognition and structures the different levels of abstraction at which intelligent agents must operate, when they interact with the environment [2]. How sub-symbolic computations transform into higher-level cognitive representations of structures is far from being understood [3]. However, there is common accepted idea that suggest that there are intermediate representations that bridge cognitive reasoning and behaviour [4]. Hence, in the design of artificial agents, a key challenge is first to find a representational space that provides an effective interface between these disparate demands, as well as a mechanism that makes proper use of this representation to interact with the environment and produce behaviour. In this sense, we approach reasoning from a behavioral perspective where the output of a reasoning process is always an action that physically interacts with the environment.
To address these challenges in a single, comprehensive system, here, we introduce a novel, brain-inspired neural network architecture that spans the domains of cognitive reasoning, perceptual inference, planning and continuous control. We follow both the emergentist and probabilistic approach to cognition, where the representational format, and the machinery to transform sensory observations into this space, is learned unsupervised but allows reasoning as an inference process. This also avoids the dependency on ground-truth labels furnished by human annotators, as these are highly labor-intensive to produce, especially when labels must cover all the relevant variables in a scene. The proposed architecture leverages the core inductive bias that the environment can be partitioned into discrete entities, or objects, that obey many useful symmetries and invariances [5]. Object-based representations and reasoning are a key tenet of perception and cognition in humans [6]. Objects constitute the environment’s separable, movable parts, as well as the logical units of reasoning and planning. They naturally take on properties of symbols, as different objects obey the same laws of physics and possess similar or analogous properties [7]. At the same time, each object representation is tethered to the sensorimotor domain via explicit attention maps. Objects, thus, are an ideal level of abstraction to interface between the disparate levels of computation that we require. We show the promise of our approach by evaluating it in tasks that require both high-level reasoning and continuous control in synthetically generated environments.
@@ 2 Related work, challenges and contribution
Structured representation learning appeared as a powerful way to introduce inductive biases and scale artificial intelligence to high-level cognition [8] and exploit symmetries and invariances that can be leveraged by decomposing scenes into their natural constituents [9]. Particularly, object-centric representations provide a natural interface between bottom-up (e.g., emergentist, connectionist) and top-down (e.g., graph based, probabilistic) approaches [7]. We summarize the closely related works that influenced the proposed architecture classified in four topics [^1] Some of the works may have overlapping features in other topics.: scene understanding (e.g., visual segmentation), physics-informed and simulation-based methods, robotics and reasoning (e.g., object relations). We analyse these works from the prism of behavioural reasoning.
Visual segmentation and tracking
Visual segmentation methods have reached maturity with high accuracy on segmentation in both static images [10] and videos [11]. From IODINE [10] that used iterative amortized inference, recent methods evolved to slot-attention architectures [12] and diffusion approaches [13]. These methods, which can track 2D and 3D objects in cluttered scenes [14], capture object collisions and discover new objects [15], are designed for perception and not for reasoning about, or controlling the segmented objects.
@@
These approaches have also shown a great potential in handling complex non-linear interactions. These are strongly influenced by the simulation approach to cognition [4]. For instance, interaction [16] and propagation [17] networks are able to understand complex scene dynamics with multiple objects and even control them to reach a desired known state. Unfortunately, these methods require full observation of the objects states. Approaches that can work directly on pixels make use of ground-truth segmentation masks [18] or Visual interaction networks [19]. While these approaches have some kind of physical “reasoning” a full-fledged architecture that performs 1st order logic reasoning from pixels and transforms to behaviour (control) is still not fully accomplished.
@@
These works focus on moving the objects meaningfully and usually have full access to the objects states (e.g. [20]). OP3 [21] is an outstanding exception. However, it has two key limitations. First, it is restricted to a discrete action space of picking up and placing objects, with no continuous control. Second, it has no ability to learn tasks—it can only plan actions towards objectives specified ad hoc by means of a goal image that shows exactly what the scene in question ought to look like. Novel approaches use ReFs to describe the 3D representation of the object and control is solved by RRTs in the latent prediction dynamics and Model Predictive Control (MPC) to execute the actions [22]. Still, ground truth objects mask are used for training and inference. Some methods have overcome this constraint through entity-based segmentation—as a simplification of object-centric representation that uses the center of the object as the location of the entity—and RL to overcome with the generation of meaningful behaviours [23]. Unfortunately, this restricts the shape of objects to “point mass” like. Besides, there are object-centric approaches that, instead of focusing on the interactive behaviour between the agent and the objects, focus on view-point matching [24] and agent navigation [25].
@@
Leaving out the literature on visual scene understanding with language, reasoning with object-centric representations as neurosymbols is, while the most promising, the least investigated area [26]. There are just a few object-centric studies on visual reasoning in static images [27]. Furthermore, there are relevant works from Large Language Models research (e.g., [28]), but they do not adhere to the grounding paradigm proposed, where reasoning should appear as an emergent property of learning a world physical model [29]. For instance in [30] the reasoning capabilities are much more expressive than our proposed approach but they use pre-trained behaviours, which are conditioned on the language generated by the LLM. Conversely, our architecture learns the world dynamics and interaction through unsupervised learning, harnessing the construction of the proto-symbols while interacting. This does not prevent the possibility of connecting proposed architecture to a LLMs similarly to [31], but through the preference network, which is already grounded.
2.1 Current challenges and architecture decisions
@@
There are two major challenges at the unsupervised learning with object-centric representations: working with complex naturalistic scene images [13] and the generation of meaningful physical actions (i.e., interacting through continuous control) derived from the reasoning process when the input is high-dimensional (i.e., image) and the environment is partially-observable. This work focus on the latter. We show that simple conditional reasoning problems, such as, “if there is a heart move squares to the left, otherwise move squares to the right” ($(A\to B)\land(\neg A\to C)$) is already unsolvable for non-object-centric representations SOTA algorithms—see Sec. 4. An underlying problem for both challenges is to achieve high performance in world state dynamic estimation (world model learning), but allowing adaptation and generalization. We designed the object-centric backbone architecture to incorporate adaptation, following the free energy principle theory of neural information processing [32, 33], where the brain estimates the world by continuously approximating its internal model, learned by experience, to the real world through approximate Bayesian computation. To implement it, we opted for an iterative variational inference approach inspired by the IODINE [10] architecture. Whilst iterative methods have shown less performance than amortized approaches (e.g., slot-attention architecture [12]), at least at perceptual tasks, it provides Bayesian filtering and smoothing, thus allowing better counterfactual reasoning. Finally, another relevant challenge, which is not addressed in this work, is to deal with complex non-linearities of physical contact, such as collisions. Physics-based simulation based approaches [18] and Deep Reinforcement Learning (RL) [23] are showing the best results but they still provide the ground truth object masks or entail oversimplified segmentation in the visual input respectively.
@@ 2.2 Contribution
This work provides a neural network architecture (Fig. 1) that can learn 1st order conditional behavioural reasoning where proto-symbols [^2] We use the term proto-symbols as in visual segmentation we use for proto-objects [34], thus describing candidates for symbols and objects. take the form of object-centic representations and perception, control and preferences (desired internal state of the agent) are dynamical processes computed through approximate Bayesian inference, in an unsupervised learning scheme from pixels. Our approach draws inspiration from natural intelligence, where the agent’s function is to generate behaviours through reasoning conditioned on perceptual cues by optimizing a single quantity: the variational free energy (or evidence lower bound) [32]. To this end, we combine bottom-up and top-down object-centric approaches [35, 36] following the emergentist approach (with unsupervised learning from pixels) but allowing probabilistic inference on the structured representation. This permits the agent to learn: $i)$ what is an object (scene understanding), $ii)$ how to mentally manipulate it through object-centric rules (reasoning) and $iii)$ generate physical behavior (i.e., continuous control). The agent can only perceive through visual input (2D image projection) of a synthetic world composed of 2D or 3D objects (see Fig. 2a for schematic). It can interact with the environment by applying forces to locations in the image, which corresponds to forces in the 2D/3D world. This architecture allows the agent to learn complex conditional rules, such as $(A\to B\land C)\land(\neg A\to D\land E)$, perform logical composition $(A\to B)\land(A\to C)\vdash A\to(B\land C)$ by learning two different rules separately but then they emerge combined during execution, and XOR operations, such as $(A\lor B)\to C)\land(A\land B\to D)$. Once the agent learns the rules, reasoning transforms into a behavioural response that changes the environment towards its internal preference. Furthermore, the agent, using the proposed architecture, shows online adaptation to unexpected changes in its environment, is robust to mild violations of its world model, and invariant to the number of objects.
@@ 3.1 Perceptual inference and generative model
To learn object-centric representations and perform inference from pixels, we use a set of $K$ weight-sharing iterative variational autoencoders (itVAEs) [37, 10], where $K$ is the number of object representations to be inferred. These itVAEs together invert a generative model in which image pixels are drawn from a Gaussian mixture distribution:
@@
where ${\bm{\mathbf{\epsilon}}}_{1_{t}}$ and ${\bm{\mathbf{\epsilon}}}_{2_{t}}$ are noise realizations drawn from a Normal distribution. Note that we are assuming that the latent representation dynamics can be captured with a 2nd-order generalized coordinates with additive noise, but this is not imposed to the environment [^3] Robotic experiments have shown that 2nd order generalized coordinates are enough to track a dynamical system [38] depending on the nature of the noise. The action ${\bm{\mathbf{a}}}_{t}^{(k)}$ on object $k$ at time $t$ is a (2-D or 3-D) vector that specifies the control command (e.g., acceleration) on the object in environment coordinates. Finally, D is a learned function that transforms the control command to its effect in the model’s latent space.
@@
Inference for each object is performed within a sliding window spanning the present and a number of past time points. Each itVAE infers beliefs for a single object and time point, taking a total of 8 inference iterations to refine these beliefs to their optimal (minimum-ELBO) values. Inference is coupled between itVAEs through information flow between represented objects and between time points in the inference window (see Appendix 6). OBR optimizes a composite ELBO loss $\mathcal{L}_{\text{comp}}=\sum_{n=1}^{N_{\text{iter}}}\frac{n}{N_{\text{iter}}}\mathcal{L}^{(n)}$ for both learning and inference, where $n$ indexes inference iterations and $\mathcal{L}$ is:

### v3

@@ Abstract
[inst2]organization=Cajal Neuroscience Center, Spanish National Research Council, city=Madrid, country=Spain
@@ 1 Introduction
A one-year old infant, before language is expressed [cowley2007human], learns the efferent-afferent patterns of sensory stimulation and motor commands to a neural representation of the environment. The pathway from the sensorium, to abstract thought, and back to the minutiae of the sensorimotor domain defines the feats of higher-order cognition and structures the different levels of abstraction at which intelligent agents must operate, when they interact with the environment [gibney2024ai]. How sub-symbolic computations transform into higher-level cognitive representations of structures is far from being understood [piantadosi2021computational]. However, there is common accepted idea that suggest that there are intermediate representations that bridge cognitive reasoning and behaviour [battaglia2013simulation]. Hence, in the design of artificial agents, a key challenge is first to find a representational space that provides an effective interface between these disparate demands, as well as a mechanism that makes proper use of this representation to interact with the environment and produce behaviour. In this sense, we approach reasoning from a behavioral perspective where the output of a reasoning process is always an action that physically interacts with the environment.
To address these challenges in a single, comprehensive system, here, we introduce a novel, brain-inspired neural network architecture that spans the domains of cognitive reasoning, perceptual inference, planning and continuous control. We follow both the emergentist and probabilistic approach to cognition, where the representational format, and the machinery to transform sensory observations into this space, is learned unsupervised but allows reasoning as an inference process. This also avoids the dependency on ground-truth labels furnished by human annotators, as these are highly labor-intensive to produce, especially when labels must cover all the relevant variables in a scene. The proposed architecture leverages the core inductive bias that the environment can be partitioned into discrete entities, or objects, that obey many useful symmetries and invariances [greff2020binding]. Object-based representations and reasoning are a key tenet of perception and cognition in humans [Peters2021]. Objects constitute the environment’s separable, movable parts, as well as the logical units of reasoning and planning. They naturally take on properties of symbols, as different objects obey the same laws of physics and possess similar or analogous properties [griffiths2010probabilistic]. At the same time, each object representation is tethered to the sensorimotor domain via explicit attention maps. Objects, thus, are an ideal level of abstraction to interface between the disparate levels of computation that we require. We show the promise of our approach by evaluating it in tasks that require both high-level reasoning and continuous control in synthetically generated environments.
@@ 2 Related work, challenges and contribution
Structured representation learning appeared as a powerful way to introduce inductive biases and scale artificial intelligence to high-level cognition [goyal2022inductive] and exploit symmetries and invariances that can be leveraged by decomposing scenes into their natural constituents [eslami2016attend]. Particularly, object-centric representations provide a natural interface between bottom-up (e.g., emergentist, connectionist) and top-down (e.g., graph based, probabilistic) approaches [griffiths2010probabilistic]. We summarize the relevant related works that influenced the proposed architecture classified in four topics [^1] Some of the works may have overlapping features in other topics.: scene understanding (e.g., visual segmentation and prediction), physics-informed and simulation-based methods, robotics and reasoning (e.g., object relations). We analyse these works from the prism of behavioural reasoning. Furthermore, Table 2 details the object-centric relevant work depending on the input and output: (A) Object control with full access to the object positions. (B) Object-centric image and video prediction and (C) Image-based object-centric control and planning.
Visual segmentation and prediction
Visual segmentation methods have reached maturity with high accuracy on image segmentation and prediction. While non-object-centric powerful video prediction methods, such as the PredRNN [wang2022predrnn], can extract object shapes and are usually used in object-based benchmarking. Here, we narrow down to object-centric solutions in both static images [Greff2019a_Iodine] and videos [elsayed2022savi++]. From IODINE [Greff2019a_Iodine] that used iterative amortized inference, recent methods evolved to slot-attention architectures [locatello2020SlotAttention] and diffusion approaches [jiang2023object]. These methods, which can track 2D and 3D objects in cluttered scenes [traub2022learning], capture object collisions and discover new objects [yu2022unsupervised], are designed for perception and not for reasoning about, or controlling the segmented objects. Connecting visual prediction algorithms to a controller usually not provide the best performance, as the features and the transition or dynamics learnt are not coupled with the action information. For instance, in world models prediction like both activated and passive pretrained options are provided [assran2025v]. Our proposed model learns a transition model in the latent space through acting in the environment, tightening the object-centric representation to the planned behaviour.
@@
These approaches have also shown a great potential in handling complex non-linear interactions. These are strongly influenced by the simulation approach to cognition [battaglia2013simulation]. For instance, interaction [battaglia2016interaction] and propagation [li2019propagation] networks are able to understand complex scene dynamics with multiple objects and even control them to reach a desired known state. Unfortunately, these methods require full observation of the objects states. Approaches that can work directly on pixels make use of ground-truth segmentation masks [piloto2022intuitive] or Visual interaction networks [watters2017visual]. While these approaches have some kind of physical “reasoning” a full-fledged architecture that performs 1st order logic reasoning from pixels and transforms to behaviour (control) is still not fully accomplished.
@@
These works focus on moving the objects meaningfully and usually have full access to the objects states (e.g. [sancaktar2022curious]). OP3 [veerapaneni2020entity] is an outstanding exception. However, it has two key limitations. First, it is restricted to a discrete action space of picking up and placing objects, with no continuous control. Second, it has no ability to learn tasks—it can only plan actions towards objectives specified ad hoc by means of a goal image that shows exactly what the scene in question ought to look like. Novel approaches use ReFs to describe the 3D representation of the object and control is solved by RRTs in the latent prediction dynamics and Model Predictive Control (MPC) to execute the actions [driess2023learning]. Still, ground truth objects mask are used for training and inference. Some methods have overcome this constraint through entity-based segmentation—as a simplification of object-centric representation that uses the center of the object as the location of the entity—and RL to overcome with the generation of meaningful behaviours [haramati2024entity]. Unfortunately, this restricts the shape of objects to “point mass” like. Besides, there are object-centric approaches that, instead of focusing on the interactive behaviour between the agent and the objects, focus on view-point matching [van2024object] and agent navigation [huo2023geovln]. In terms of architecture, according to the classification provided in [garrido2024learning], our approach would lie into the JEPA models, where prediction happens in the latent space and includes explicit control. While V-JEPA 2-AC [assran2025v] uses a more powerful embedding encoding than our dynamic iterative variational inference perceptual backbone, they still suffer the same limitation as in other goal-image-conditioned manipulation: “the optimization target assumes that we have access to visual goals”. We address this limitation in OBR by including a preference network, allowing behavioural 1st order logic. Another important factor is the inference time. [assran2025v] reports an inference time of 16 seconds per action plan (using cross-entropy optimization), and this is not still suitable for a real-world setting. While our current architecture does not generalize to realistic images, action computation is computed in closed form, achieving fast inference times (less than one second). Recent architectures combining slot-attention and model-based RL [sold2025mosbach] are also providing suitable inference times but their training time is comparably much higher.
@@
Leaving out the literature on visual scene understanding with language, reasoning with object-centric representations as neurosymbols is, while the most promising, the least investigated area [assouel2022object]. There are just a few object-centric studies on visual reasoning in static images [webb2024systematic]. Furthermore, there are relevant works from Large Language Models research (e.g., [driess2023palm]), but they do not adhere to the grounding paradigm proposed, where reasoning should appear as an emergent property of learning a world physical model [taniguchi2023world]. For instance in [ahn2022can] the reasoning capabilities are much more expressive than our proposed approach but they use pre-trained behaviours, which are conditioned on the language generated by the LLM. Conversely, our architecture learns the world dynamics and interaction through unsupervised learning, harnessing the construction of the proto-symbols while interacting. This does not prevent the possibility of connecting proposed architecture to an LLM similarly to [xu2023llms], but through the preference network, which is already grounded.
Table 1: Relevant approaches related to object centric behavior generation. We categorized the methods into three blocks: (A) Object control with full access to the object positions, (B) Object-based image and video prediction—while they do not encode actions, they provide the architectures for high-dimensional object-centric control solutions—and (C) Image-based object-centric control and planning
Methods Input/Action Perception Transition Control Goal input Objective Strengths / Limitations
[…]
Interaction Net [battaglia2016interaction] Obj. pos / - - graph encoding No - - Predicts non-linear interactions (visual encoder version exists).
Propagation Net[li2019propagation] Obj. pos / Position or velocity - graph encoding SGD +MLP Desired obj. positions Chanfer distance Allows control of complex non-linear interactions. / needs ground truth object positions.
CEE-US [sancaktar2022curious] Obj. pos / End-effector pos (2D plane) - Graph Neural Networks RL Desired obj. positions Max E[r] Incorporates curiosity drivers. / needs ground truth object positions.
[…]
IODINE [Greff2019a_Iodine] Image/- IAI [Marino2018] No - - - Perception works like a filter with the refinement network. / Only static images.
SlotAttention [locatello2020SlotAttention] Image/- Slot Attention GRU - - - Reduced inference time for object representation.
SAVI [kipfconditional_SAVi, elsayed2022savi++] Video/- Slot Attention GRU - - - Works for realistic videos / Tends to smooth the shapes into blobs and does not incorporate control.
LOCI [traub2023learning] Video/- Autoencoder Gatel0rd[gumbsch2021sparsely] - - - Uses explicit 2D positional encodings as Gaussians.
[…]
OP3 [veerapaneni2020entity] Img seq. / Discrete (primitives) IAI[Marino2018] Approximate dynamics network Cross-Entropy Optimization Encoded goal image Min latent distance. Coupled action and perception / only allows discrete actions
SMORL [zadaianchuk2020self] Img seq. / End-effector pos (2D plane) SCALOR[jiang2019scalor] Attention head SAC+HER Encoded goal image Distance between object position and goal if they are similar in features Allows RL framework integration / Scales poorly to the number of objects.
Entity RL[veerapaneni2020entity] Img seq./ EE joint position Deep Latent Particles [daniel2022unsupervised] Transformer TD3+HER Reward (distance between state and goal particle). Max E[r] Scales linearly with the number of objects / cannot encode complex object properties (e.g., shapes).
SOLD [sold2025mosbach] Image seq./ Continuous positions SAVI [kipfconditional_SAVi] Transformer [villar2023object] Dreamer-like RL Task reward Max E[r] + Entropy regularizer Beats monolitic approaches in relational reasoning / Very costly to train and shapes are not properly captured.
OBR (proposed) Img seq. / Continuous (3D acc. vectors) IAI [Marino2018] Latent linear dynamics Closed-form Expected Free Energy [millidge2021whence] Preference network (set-structured MLP) composite ELBO loss Allows 1st order proto-object behavioural reasoning and fast action computation. / Objects’ interactions are not explicitly modelled.
2.1 Current challenges
@@
There are three major challenges in unsupervised learning with object-centric representations: $i)$ working with complex naturalistic scene images [jiang2023object], ii)ii) the generation of meaningful physical actions (i.e., interacting through continuous control) when the input is high-dimensional (i.e., image) and the environment is partially observable and $iii)$ behavioural reasoning, where these actions adhere to a high-order goal, such as logics. This work focuses on the latter. We show that simple conditional reasoning problems, such as, “if there is a heart move squares to the left, otherwise move squares to the right” ($(A\to B)\land(\neg A\to C)$) is already unsolvable for non-object-centric representations SOTA algorithms—see Sec. 4. An underlying problem for the challenges is to achieve high performance in world state dynamic estimation (world model learning), but allowing adaptation and generalization. We designed the object-centric backbone architecture to incorporate adaptation, following the free energy principle theory of neural information processing [friston2010free, lanillos2021active], where the brain estimates the world by continuously approximating its internal model, learned by experience, to the real world through approximate Bayesian computation.
Finally, another relevant challenge, which is not addressed in this work, is to deal with complex non-linearities of physical contact, such as collisions. Physics-based simulation-based approaches [piloto2022intuitive] and Deep Reinforcement Learning (RL) [haramati2024entity, sold2025mosbach] variants are showing the best results.
@@ 2.2 Contribution
This work provides a neural network architecture (Fig. 1) that can learn 1st order conditional behavioural reasoning where proto-symbols [^2] We use the term proto-symbols as in visual segmentation we use for proto-objects [walther2006modeling], thus describing candidates for symbols and objects. take the form of object-centic representations and perception, control and preferences (desired internal state of the agent) are dynamical processes computed through approximate Bayesian inference, in an unsupervised learning scheme from pixels. Our approach draws inspiration from natural intelligence, where the agent’s function is to generate behaviours through reasoning conditioned on perceptual cues by optimizing a single quantity: the variational free energy (or evidence lower bound) [friston2010free]. To this end, we combine bottom-up and top-down object-centric approaches [rasouli2020attention, van2023object] following the emergentist approach (with unsupervised learning from pixels) but allowing probabilistic inference on the structured representation. This permits the agent to learn: $i)$ what is an object (scene understanding), ii)ii) how to mentally manipulate it through object-centric rules (reasoning) and $iii)$ generate physical behavior (i.e., continuous control). The agent can only perceive through visual input (2D image projection) of a synthetic world composed of 2D or 3D objects (see Fig. 2a for schematic). It can interact with the environment by applying forces to locations in the image, which corresponds to forces in the 2D/3D world. This architecture allows the agent to learn complex conditional rules, such as $(A\to B\land C)\land(\neg A\to D\land E)$, perform logical composition $(A\to B)\land(A\to C)\vdash A\to(B\land C)$ by learning two different rules separately but then they emerge combined during execution, and XOR operations, such as $(A\lor B)\to C)\land(A\land B\to D)$. Once the agent learns the rules, reasoning transforms into a behavioural response that changes the environment towards its internal preference. Furthermore, the agent, using the proposed architecture, shows online adaptation to unexpected changes in its environment, is robust to mild violations of its world model, and invariant to the number of objects.
@@ 3.1 Perceptual inference and generative model
To learn object-centric representations and perform inference from pixels, we use a set of $K$ weight-sharing iterative variational autoencoders (itVAEs) [Marino2018, Greff2019a_Iodine], where $K$ is the number of object representations to be inferred. These itVAEs together invert a generative model in which image pixels are drawn from a Gaussian mixture distribution:
@@
where ${\bm{\mathbf{\epsilon}}}_{1_{t}}$ and ${\bm{\mathbf{\epsilon}}}_{2_{t}}$ are noise realizations drawn from a Normal distribution. Note that we are assuming that the latent representation dynamics can be captured with a 2nd-order generalized coordinates with additive noise, but this is not imposed to the environment [^3] Robotic experiments showed that 2nd order generalized coordinates are enough to track a dynamical system [bos2022free] depending on the noise source. The action ${\bm{\mathbf{a}}}_{t}^{(k)}$ on object $k$ at time $t$ is a (2-D or 3-D) vector that specifies the control command (e.g., acceleration) on the object in environment coordinates. Finally, D is a learned function that transforms the control command to its effect in the model’s latent space.
@@
Inference for each object is performed within a sliding window spanning the present and a number of past time points.
Each itVAE infers beliefs for a single object and time point, taking a total of 8 inference iterations to refine these beliefs to their optimal (minimum-ELBO) values. Inference is coupled between itVAEs through information flow between represented objects and between time points in the inference window (see Appendix 6). OBR optimizes a composite ELBO loss $\mathcal{L}_{\text{comp}}=\sum_{n=1}^{N_{\text{iter}}}\frac{n}{N_{\text{iter}}}\mathcal{L}^{(n)}$ for both learning and inference, where $n$ indexes inference iterations and $\mathcal{L}$ is:
@@
where ${\bm{\mathbf{\Psi}}}_{t}$ defines the action field— described in the next section— that maps object-centric actions into to real actions.
Action mapping module
A key challenge in multi-object environments is that an agent’s internal representation of objects has an unknown (and possibly imperfect) correspondence to the objects in the environment. Even if the object features are inferred with perfect accuracy, the order of the objects in the representation is arbitrary. To solve this correspondence problem, we define an action space (e.g., accelerations) that can be placed at pixel locations within the environment’s image grid, and objects receive the sum of all accelerations that coincide with their visible pixels. Specifically, we introduce the notion of an action field, which transform into real actions if the pixel belongs to a specific object in the environment, ${\bm{\mathbf{\Psi}}}=[{\bm{\mathbf{\psi}}}_{1},...,{\bm{\mathbf{\psi}}}_{M}]^{T}$: an $M\times 2]$ matrix (with $M$ the number of pixels in an image or video frame), such that the $i$-th row in this matrix (${\bm{\mathbf{\psi}}}_{i}$) specifies the (x,y,z)-acceleration applied at pixel $i$. The action on the $k$-th object is then given by:
${\bm{\mathbf{a}}}_{t}^{(k)}=\sum_{i}[m_{i}=k]{\bm{\mathbf{\psi}}}_{i}$ (5)
The action inference module does not incorporate a decoder network, as the quality of the action beliefs is computed by evaluating equation 5 and plugging this into the ELBO loss from equation 4. While this requires some additional sampling operations (see Appendix 7), no neural network is required for this. This module does include a (shallow) refinement network—detailed in the Appendix 6. This network takes as input the current variational parameters ${\bm{\mathbf{\lambda_{a}}}}^{(k)}$ (means and variances), their gradients, and the ‘expected object action’, $\sum_{i}\hat{m}_{ik}{\bm{\mathbf{\psi}}}_{i}$.